Show that the Mercator projection
σ(u ,v) = (sech u cos v, sech u sin v ,tanh u)
is a regular surface patch of the unit sphere
1
Expert's answer
2020-10-19T18:16:59-0400

Jacobian= "\begin{bmatrix}
 \partial f_{1}/\partial u & \partial f_{1}/\partial v \\
 \partial f_{2}/\partial u & \partial f_{2}/\partial v\\
 \partial f_{3}/\partial u&\partial f_{3}/\partial v 
\end{bmatrix}" = "\begin{bmatrix}
 -(sech \ u \ tanh \ u)\cos v & -sech \ u \ \sin v\\ 
 -(sech \ u \ tanh \ u)\sin v & sech \ u \ \cos v\\ 
sech^2 \ u &0 
\end{bmatrix}" . Now if we look at the minor deleting the first row we get "sech^3 \ u \cos v." Now "sech \ u\neq 0" , hence minor vanishes if "\cos v=0." Now taking the minor after deleting 2nd row gives, "-sech^3 \ u \sin v." This shows "\sin v=0." But both "\sin v , \cos v" cannot be zero hence one of the two minors never vanish showing the Jacobian has rank two. Also, all points lie on the sphere since "(sech \ u)^2 \cos^2u+(sech \ u)^2\sin^2u+(tan \ h)^2u=1."